# The generative quantum eigensolver (GQE) and its application for ground state search

### v1

@@ 2.2 GPT Quantum Eigensolver
where $\mathcal{Z}=\sum_{\vec{j}}\exp\left(-\beta w_{\textrm{sum}}(\vec{j})\right)$ and we write the hyper-parameter $\beta$ explicitly.
Training
To construct the training scheme for GPT-QE, let us consider the process sampling $U_{N}(\vec{j})$ according to $p_{N}(\beta,\vec{j})$ and applying it to $\rho_{0}$. Let $\rho(\beta)$ be the quantum state generated by the stochastic process. With $\mathcal{E}_{N}(\vec{j},\rho):=U_{N}(\vec{j})\rho U_{N}(\vec{j})^{\dagger}$ , it can be written as
We observe that if $w_{\textrm{sum}}(\vec{j})=E_{N}(\vec{j})$ is satisfied, $\rho(\beta)$ gives a pseudo thermal state with the inverse temperature $\beta$ in the sense that the quantum state is generated according to the probability $\exp\left(-\beta E_{N}(\vec{j})\right)$. Therefore, increasing the value of $\beta$ creates a bias towards generating lower energy quantum states. We note that $\rho(\beta)$ is not, in general, the exact thermal state since the quantum state $\mathcal{E}_{N}(\vec{j},\rho_{0})$ cannot represent all the eigenstates in the whole Hilbert when $N$ and $L$ are constrained.
From this observation, we design our scheme for the training/pre-training so that $w_{\textrm{sum}}(\vec{j})\simeq E_{N}(\vec{j})$ is satisfied. More specifically, in each iteration, we sample $\{\vec{j}_{m}\}_{m=1}^{M}$, calculate the cost function
$\begin{split}&C\left(\{w_{\textrm{sum}}(\vec{j}_{m})\}_{m=1}^{M},\{E_{N}(\vec{j}_{m})\}_{m=1}^{M}\right)\\ &=\frac{1}{M}\sum_{m=1}^{M}\tilde{C}\left(w_{\textrm{sum}}(\vec{j}_{m}),E_{N}(\vec{j}_{m})\right),\\ &\tilde{C}(w_{\textrm{sum}}(\vec{j}),E_{N}(\vec{j})):=\left(e^{-w_{\textrm{sum}}(\vec{j})}-e^{-E_{N}(\vec{j})}\right)^{2},\end{split}$ (5)
and update the parameters in GPT by backpropagation. Since we match the sum of logits with the energy function, we call this technique logit-matching. We overview the training process in Fig. 3. It should be noted that $M$, which corresponds to the batch size of the data in the context of the machine learning, is a hyper-parameter we can choose, and it does not explicitly depend on $N$ and $L$. Therefore, in principle, we can freely choose the number of quantum circuit runs in each iteration. However, we must also be aware that small $M$ leads to a significant statistical error in the cost function estimation. The best strategy to choose $M$ should be studied in future work. It is also possible to optimize $M$ by using hyper-parameter optimization, which itself is an active area of exploration in the field of machine learning. We also note that the parameter $\beta$ can be used to control the trade-off between exploitation and exploration. In other words, by adjusting $\beta$, we can influence how the algorithm balances between intensively searching within a particular area (exploitation) and exploring new, potentially promising areas (exploration). A higher value of $\beta$ typically encourages more exploitation, focusing the search on areas already identified as having high potential. Conversely, a lower $\beta$ value promotes exploration, allowing the algorithm to investigate a broader range of solutions that may lead to discovering better-performing options not yet explored. In our numerical experiment, we initially set $\beta$ to a small value and then gradually increase it over time.
Pre-Training
By applying the logit-matching technique, we can easily incorporate a pre-training scheme in GPT-QE (see Fig. 3). Suppose we have the dataset $\mathcal{D}=\{\vec{j}_{m},E_{N}(\vec{j}_{m})\}_{m=1}^{M_{D}}$. In pre-training, we input each $\vec{j}_{m}$ to GPT and obtain $w_{\textrm{sum}}(\vec{j}_{m})$. Then, we update parameters in GPT so that $w_{\textrm{sum}}(\vec{j}_{m})$ becomes close to $E_{N}(\vec{j}_{m})$ from the dataset by using the cost function (5). We note that we can split the dataset $\mathcal{D}$ into $B$ batches $\{\mathcal{D}_{b}\}_{b=1}^{B}$ and train GPT with each batch.
The pre-training process is entirely classical, eliminating the need to use quantum devices as long as datasets are available. The question then arises: How do we obtain these datasets? Primarily, they can be sourced from previous GPT-QE training. Specifically, previous quantum evaluations performed as part of solving related tasks can be used to construct such a dataset. This opens up the possibility of a large-scale effort to simulate several Hamiltonians (e.g. molecules and materials) in the cloud using high-performance quantum simulators or on actual quantum devices and gaining performance for GQE as the pertaining dataset becomes more comprehensive over time.
We propose three scenarios to leverage the pre-training scheme effectively: (i) model-to-model transfer, (ii) config-to-config transfer, and (iii) molecule-to-molecule transfer. Below, we describe these scenarios in detail.
(i) Model-to-model transfer scenario
In the model-to-model transfer scenario, we assume that the Hamiltonian used for data generation and the one we want to get the ground state of are the same. We first prepare a model represented as $\texttt{GPT}_{A}$ and train it by using the Hamiltonian $\hat{H}$. While training $\texttt{GPT}_{A}$, we obtain $\mathcal{D}=\{\vec{j}_{m},E_{N}(\vec{j}_{m})\}_{m=1}^{M}$. The result of the training may be unsatisfactory, and we try another model, represented as $\texttt{GPT}_{B}$. For example, $\texttt{GPT}_{B}$ may have more attention layers than $\texttt{GPT}_{A}$. Then, the dataset $\mathcal{D}$ can be utilized for the pre-training of $\texttt{GPT}_{B}$ and we can then run the training algorithm with $\texttt{GPT}_{B}$ after the pre-training. It is also possible that the dataset $\mathcal{D}$ is used for the training of $\texttt{GPT}_{A}$ itself to obtain a better initialization (Experience replay).
@@
Above, we describe the case where GPT-QE obtains the dataset. However, it is also possible that the dataset could be generated from other algorithms, e.g., tensor network calculations and VQE, if we can (approximately) convert the data obtained in those algorithms to a data format acceptable for GPT-QE.
(ii) Config-to-config transfer scenario
In the config-to-config transfer scenario, we utilize the data obtained in training to find the ground state of a Hamiltonian $\hat{H}$ corresponding to one spatial nuclear configuration to that of another Hamiltonian $\hat{H}^{\prime}$ corresponding to a different nuclear configuration. We still assume that $\hat{H}$ and $\hat{H}^{\prime}$ are Hamiltonians of the same molecule, albeit at different geometries. In the context of VQA, similar approaches have already been proposed [31, 32]. Let us first assume that the Hamiltonian can be parameterized as follows:
$\hat{H}(\vec{\Delta})=\sum_{a=1}^{N_{H}}h_{a}(\vec{\Delta})\hat{P}_{a},$ (6)
where $\vec{\Delta}$ is a set of parameters corresponding to a configuration, $N_{H}$ is the number of terms in the Hamiltonian, each $h_{a}(\vec{\Delta})$ is a real-valued function of the parameters $\vec{\Delta}$, and each $\hat{P}_{a}$ is a tensor product of Pauli operators. An example of the parameters in $\vec{\Delta}$ is the bond length; when we change the bond length, the set of $\{\hat{P}_{a}\}_{a=1}^{N_{H}}$ does not change, and only the coefficients in the Hamiltonian change. With this parameterization scheme, we can transfer the dataset obtained in the training of $\hat{H}(\vec{\Delta})$ to a dataset available for the pre-training with the Hamiltonian $\hat{H}(\vec{\Delta}^{\prime})$.
We propose two methods to realize the config-to-config transfer, which can be effectively combined with each other. One method is the following coefficient re-weighting (see also [31]). In the training with $\hat{H}(\vec{\Delta})$, we estimate $\{q_{a}(\vec{j})\}_{a=1}^{N_{H}}$ defined by $q_{a}(\vec{j}):=\textrm{Tr}\left(\hat{P}_{a}\mathcal{E}_{N}(\vec{j},\rho_{0})\right)$ and combine them to estimate the energy as
$E_{N}\left(\vec{\Delta},\vec{j}\right)=\sum_{a=1}^{N_{H}}h_{a}(\vec{\Delta})q_{a}(\vec{j}),$ (7)
where we simply write $E_{N}\left(\hat{H}(\vec{\Delta}),\vec{j}\right)$ as $E_{N}\left(\vec{\Delta},\vec{j}\right)$.
The estimated values $\{q_{a}(\vec{j})\}_{a=1}^{N_{H}}$ are also usable to construct the estimation value of $E_{N}(\vec{j},\vec{\Delta}^{\prime})$ with a different configuration, i.e., we can simply combine the measured Pauli expecectation values using different coefficients $h_{a}(\vec{\Delta}^{\prime})$. By this process, we obtain the dataset $\mathcal{D}=\left\{\vec{j}_{m},E_{N}\left(\vec{\Delta}^{\prime},\vec{j}_{m}\right)\right\}_{m=1}^{M}$, which can then be used for the pre-training when solving the search for the ground state of the Hamiltonian $H(\vec{\Delta}^{\prime})$.
Another method to achieve config-to-config transfer is adding extra inputs $\vec{\Delta}$ to the GPT function. More specifically, we would imagine extending GPT so that it takes the configuration $\vec{\Delta}$ as its input in addition to the variable inputs $\{j_{1},\cdots j_{k}\}$ and outputs the logits $W^{(k)}$. Let us write the extended function as $\texttt{GPT}_{+}$.
By training $\texttt{GPT}_{+}$ with the different configurations $\vec{\Delta}_{1},\cdots,\vec{\Delta}_{R}$, we obtain the datasets $\mathcal{D}$ = $\{\mathcal{D}_{r}\}_{r=1}^{R}$, where $\mathcal{D}_{r}:=\{\vec{\Delta}_{r},\vec{j}_{m_{r}},E_{N}(\vec{\Delta}_{r},\vec{j}_{m_{r}})\}_{m_{r}=1}^{M_{r}}$ and each $M_{r}$ is the number of sequences generated in the training with each configuration. Then, the dataset $\mathcal{D}$ can be used for the pre-training when we search for the ground state of $\hat{H}(\vec{\Delta}^{\prime})$ of a new configuration $\vec{\Delta}^{\prime}$. In the pre-training, $\vec{\Delta}_{r}$ is used as inputs of $\texttt{GPT}_{+}$ as well as $\vec{j}_{m_{r}}$, where $\vec{\Delta}^{\prime}$ is used as inputs in the main training.
(iii) Molecule-to-molecule transfer scenario
In the molecule-to-molecule transfer scenario, we propose utilizing the dataset generated from the training of one molecule for the pre-training of another. Implementing this approach requires a significant extension of the model. At this stage, we focus on outlining the concept rather than delving into the specifics of the implementation. In this scenario, it is crucial to use the information of the molecule, such as the Hamiltonian, as an input, similar to our approach in the $\texttt{GPT}_{+}$ model for config-to-config transfer.
Figure 4: The training results with the electronic structure Hamiltonians: $\texttt{H}_{2}$ (top left), LiH (top right), $\texttt{BeH}_{2}$ (bottom left), and $\texttt{N}_{2}$ (bottom right) in sto-3g basis. The results of GPT-QE are depicted as green points (gpt-qe). The Hartree-Fock energy is indicated by a gray dotted line (hf), and the exact full configuration interaction energy, calculated through diagonalization, is represented by a black line (exact). For GPT-QE, the best result in three trials is shown.
Figure 5: The corresponding circuit for the GPT-QE result of the ${\texttt{H}}_{2}$ Hamiltonian at a bond length of 2.0 angstroms. The initial two X gates serve to prepare the Hartree-Fock state. The circuit is depicted across two lines for clarity. Each number positioned above a circuit component indicates its respective token number. Although there are a total of 10 tokens, one of these corresponds to the identity operator and is not shown in the diagram.
Figure 6: Training performance with the $\texttt{N}_{2}$ Hamiltonian, comparing scenarios with and without pre-training. The left figure shows the minimum energy found by the step during the trial, while the right figure illustrates the mean energy of the quantum states generated at each step. The results from the training with pre-training are indicated by a green line, and those from the training without pre-training are shown by a gray line. In both figures, the mean and its standard deviation across ten random initializations is drawn.
@@
The primary challenge lies in uniformly treating molecules with varying numbers and types of molecular orbitals. For instance, there is not a straightforward mapping between the molecular orbitals of $\texttt{H}_{2}$ and those of $\texttt{H}_{2}\texttt{O}$ as determined by Hartree-Fock calculations. Therefore, the gate sequence in the calculation of $\texttt{H}_{2}$ does not have an a priori correspondence with that in $\texttt{H}_{2}\texttt{O}$. Finding transferable representations of electronic structure information that are useful for machine learning applications is a vibrant research field on its own [33, 34]. One potential solution to this challenge in our setting is to write the molecular Hamiltonian employing the basis of atomic orbitals instead of that of molecular orbitals. More specifically, we define the creation and the annihilation operators for each atomic orbital and correspond them with qubits. Then, a gate sequence for $\texttt{H}_{2}$ has an analogous physical meaning when applied to the portion corresponding to hydrogen atoms in $\texttt{H}_{2}\texttt{O}$. This approach facilitates effective molecule-to-molecule transfer, particularly between molecules with the same atomic composition. Alternatively, one could devise a quantum circuit that couples molecular fragments encoded initially in independent qubits. In this manner, the optimized gates for the fragments can be invoked as a starting point for the full molecular circuit.
@@ 3 Results
In this section, we showcase the effectiveness of training and pre-training in GPT-QE for approximating ground states using electronic structure Hamiltonians. We use the molecular Hamiltonians of $\texttt{H}_{2}$, LiH, $\texttt{BeH}_{2}$, and $\texttt{N}_{2}$ in the sto-3g basis for this purpose.
The configuration of the GPT-QE model is as follows. Our operator pool is a set of Pauli time evolutions: $\mathcal{G}=\{e^{i\hat{P}_{j}t_{j}}\}_{j}$, where $\hat{P}_{j}$ represents a tensor product of Pauli operators, and $t_{j}$ is a real value. Defining $\mathcal{P}$ as the set of $\hat{P}_{j}$ and $\mathcal{T}$ as the set of $t_{j}$, the size of the operator pool is $|\mathcal{P}|\times|\mathcal{T}|$. For $\mathcal{P}$, we derive chemically inspired choices from the unitary coupled-cluster single and double excitations (UCCSD). Letting $T$ denote the sum of all fermionic excitation operators included in UCCSD, $\mathcal{P}$ is selected such that $e^{iP_{\ell}\theta_{\ell}}(P_{\ell}\in\mathcal{P})$, with an angle $\theta_{\ell}$,is part of the decomposed operators when $e^{T-T^{\dagger}}$ is broken down into Pauli time evolutions by the Trotter decomposition [35]. The identity operator is also included in $\mathcal{P}$. For $\mathcal{T}$, we choose $\mathcal{T}=\left\{\pm 2^{k}/160\right\}_{k=1}^{4}$. Regarding the transformer model, we employ a configuration identical to that of GPT-2 [7], featuring 12 attention layers, 12 attention heads, and 768 embedding dimensions. The initial state, $\rho_{\textrm{0}}$, is set to the Hartree-Fock state. For numerical stability, we add an offset to the output of the quantum device. More specifically, when calculating the cost function (5), we substitute $E_{N}(\vec{j})+E_{\textrm{offset}}$ instead of the original $E_{N}(\vec{j})$. The value of $E_{\textrm{offset}}$ is chosen to be 00, $7$, $14$, and $106$ for $\texttt{H}_{2}$, LiH, $\texttt{BeH}_{2}$, and $\texttt{N}_{2}$, respectively.
@@ 3.1 Training Performance
In Fig. 4, we present the results of GPT-QE training for each electronic structure Hamiltonian. The horizontal axis represents the bond length, while the vertical axis shows the energy value. The results of GPT-QE are depicted as green points (gpt-qe). The Hartree-Fock energy is indicated by a gray dotted line (hf), and the exact full configuration interaction energy, calculated through diagonalization of the full Hamiltonian, is represented by a black line (exact).
For $\texttt{H}_{2}$, we run 50 steps with a number of samples per step set at $M=25$. And for the other molecules, we run 500 steps with $M=50$. The inverse temperature $\beta$ starts at $5$ and is increased by $0.1$ at each step. The number of tokens, i.e., the length of the output sequence, is fixed at 10 for $\texttt{H}_{2}$ and 40 for the other molecules. We conducted three trials of GPT-QE and recorded the minimum energy in each. The plotted points represent the best results from these trials.
We observe that GPT-QE effectively identifies low-energy states closely approximating the ground state. In Appendix A, we also compare the average performance of GPT-QE with the case when quantum gates are randomly generated from the operator pool $\mathcal{G}$ and verify the effectiveness of the GPT-QE’s training scheme, which employs logit-matching to steer the transformer towards desired outcomes. Some of the results include errors larger than the chemical accuracy. Such errors can be further reduced by choosing suitable hyper-parameters and pre-training, as described in the upcoming subsection. It should also be noted that we can enhance the accuracy by applying VQE as a post-processing step.
For demonstration purposes, Fig 5 displays the quantum circuit corresponding to the data point for a bond length of 2.0 Å, which is drawn using an open-source quantum circuit visualization tool [37]. To our knowledge, this is the first instance of a transformer-generated quantum circuit reported in a scientific publication.

### v2

@@ 2.2.1 Quantum circuits generation in GPT-QE
where $\mathcal{Z}$ is a normaliziation factor and we write the hyper-parameter $\beta$ explicitly.
2.2.2 Training
The training process in GPT-QE involves two main phases: data collection through circuit sampling and model updates using batched data from the replay buffer. The replay buffer is a technique commonly used in reinforcement learning [25] to store and reuse past experiences for more efficient learning. We implement training with a replay buffer of fixed size $N_{\text{buffer}}^{\text{max}}$ that operates as a first-in-first-out (FIFO) queue to store and manage previously sampled circuits and their corresponding energies.
In each training epoch, we collect $N_{\text{sample}}$ new circuit samples through the process described in section 2.2.1. We evaluate the corresponding energies of circuits using the quantum device (or its simulator). These sequence-energy pairs are then stored in the replay buffer using a FIFO mechanism. The replay buffer contains a dataset $\mathcal{Q}=\{\vec{j}_{m},E(\vec{j}_{m})\}_{m=1}^{N_{\text{buffer}}}$ where $N_{\text{buffer}}$ represents the current number of entries in the buffer. When the buffer reaches its maximum capacity $N_{\text{buffer}}^{\text{max}}$, the oldest entries are automatically removed to make room for new data. Note that when $N_{\text{buffer}}^{\textrm{max}}=N_{\text{sample}}$, this means the newly collected data is used immediately and then disposed of.
The model update phase utilizes data from the replay buffer to improve the model’s performance. During this phase, the $N_{\text{buffer}}$ data points in the replay buffer are processed in batches of size $N_{\text{batch}}$, and this process is repeated for $N_{\text{iter}}$ iterations. In each iteration, the loss function is computed using the current batch of data points, and the model parameters are updated through gradient descent to minimize the loss. The loss function is designed so that logits corresponding to high-energy circuits become smaller, while logits corresponding to low-energy circuits become larger. This encourages the model to generate circuits with lower energies. As our specific construction of the loss functions, we use the logit matching loss and Group Relative Policy Optimization (GRPO) [26] loss in our experiment, as introduced in Appendix B.
The nature of the training in the GPT-QE approach naturally enables pre-training capabilities. By storing previous learning experiences in the replay buffer, we can leverage this accumulated data to pre-train models for new tasks or configurations, providing a foundation for learning across different quantum optimization problems. The pre-training process is entirely classical, eliminating the need to use quantum devices as long as datasets are available.
In order to reuse the data in the replay buffer for a different problem, the energies of the replay buffer should be translated to the energies of the new Hamiltonian $H$. For this purpose, we utilize that, when computing on quantum devices, we evaluate the expectation values for each Pauli basis, from which the expectation value of the target Hamiltonian is constructed. These individual Pauli expectation values can be stored and reused for estimating different observables, significantly improving computational efficiency. In particular, it allows the energy of $H^{\prime}$ to be estimated if it contains the same Pauli strings as the original Hamiltonian. The most prominent example of this approach is the electronic structure Hamiltonian, where the same set of Pauli expectation values can be used to estimate energies for different molecular geometries. This technique is detailed in Appendix C and investigated in Section 3.2 (see also [27]).
@@ 2.2.2 Training
Above, we describe the case where GPT-QE itself constructs the dataset. However, it is also possible that the dataset could be generated from other algorithms, e.g., tensor network calculations and VQE, if we can (approximately) convert the data obtained in those algorithms to a data format acceptable for GPT-QE. Following our initial submission, subsequent work has demonstrated this approach in quantum combinatorial optimization [28].
2.2.3 Conditional generalization
A key advantage of using generative models is the ability to incorporate conditional inputs, enabling the generation of problem-specific quantum circuits based on classical information such as molecular geometry or optimization constraints. We can define a conditional GPT model as $\texttt{GPT}(\vec{\Delta})$, where $\vec{\Delta}$ represents the conditional input parameters (e.g., molecular geometry, bond lengths, or other problem-specific constraints). This model can generate quantum circuits that are tailored to specific parameter values $\vec{\Delta}$, and the corresponding cost can be evaluated using the Hamiltonian $\hat{H}(\vec{\Delta})$ for that particular configuration. When a new parameter configuration $\vec{\Delta}_{\text{new}}$ is encountered, full retraining is likely not required—the model can immediately generate appropriate quantum circuits for the new configuration using the learned conditional mapping, though fine-tuning remains an option for further optimization. This capability opens up new possibilities for generalization across electronic structure problems, where we can leverage knowledge gained from one system to improve performance on related systems. While this paper does not provide actual demonstrations of these generalization capabilities, we outline the promising directions and potential applications.
One approach is config-to-config transfer, where we train the model on a set of molecular geometries and then apply it to new geometries of the same molecule. This would extend the pre-training framework by utlilizing a single model across multiple geometries, rather than re-using data for training a separate model for each geometry. Another promising direction is molecule-to-molecule transfer, where we leverage datasets generated from training on one molecule to improve performance on chemically related molecules. As mentioned, both approaches require extending the model to incorporate molecular information as conditional inputs. While this represents a more challenging extension, it could enable the development of general-purpose quantum circuit generators that can adapt to different molecular systems.
@@ 2.2.3 Conditional generalization
The primary challenge for the molecule-to-molecule transfer lies in uniformly treating molecules with varying numbers and types of molecular orbitals. For instance, there is not a straightforward mapping between the molecular orbitals of $\texttt{H}_{2}$ and those of $\texttt{H}_{2}\texttt{O}$ as determined by Hartree-Fock calculations. Therefore, the gate sequence in the calculation of $\texttt{H}_{2}$ does not have an a priori correspondence with that in $\texttt{H}_{2}\texttt{O}$. Finding transferable representations of electronic structure information that are useful for machine learning applications is a vibrant research field on its own [29, 30]. One potential solution to this challenge in our setting is to write the molecular Hamiltonian using the basis of atomic orbitals instead of that of molecular orbitals. More specifically, we define the creation and the annihilation operators for each atomic orbital and correspond them with qubits. Then, a gate sequence for $\texttt{H}_{2}$ has an analogous physical meaning when applied to the portion corresponding to hydrogen atoms in $\texttt{H}_{2}\texttt{O}$. This approach may facilitate effective molecule-to-molecule transfer, particularly between molecules with the same atomic composition. Alternatively, one could devise a quantum circuit that couples molecular fragments encoded initially in independent qubits. In this manner, the optimized gates for the fragments can be invoked as a starting point for the full molecular circuit.
Note that following our initial preprint, subsequent work has demonstrated the effectiveness of conditional generalization approaches in quantum combinatorial optimization, where conditional inputs have been successfully used to generate problem-specific quantum circuits [18, 28]. These developments validate the potential of our generative approach for conditional generalization across different quantum optimization problems.
(a) $\texttt{H}_{2}$ (4 qubits)
(b) LiH (10 qubits)
(c) $\texttt{BeH}_{2}$ (12 qubits)
(d) $\texttt{N}_{2}$ (12 qubits)
Figure 4: Training results for each molecule showing energy vs. bond length plots with GQE results (green points), HF energy (gray dotted line), CCSD energy (blue dashed line), and FCI energy (black line). The number of tokens (circuit length) is set to 10 for $\texttt{H}_{2}$, 40 for LiH, 60 for $\texttt{BeH}_{2}$, and 100 for $\texttt{N}_{2}$. Note that there is a jump in the FCI calculation for $\texttt{N}_{2}$ around 0.93-0.94 Å, which is due to a change in the active space selection in the CAS calculation. This does not affect the objectives or conclusions of this experiment since it affect all approaches equally.
(a) $\texttt{H}_{2}$ (4 qubits)
(b) LiH (10 qubits)
(c) $\texttt{BeH}_{2}$ (12 qubits)
(d) $\texttt{N}_{2}$ (12 qubits)
Figure 5: Absolute error from FCI on a logarithmic scale for each molecule, with chemical accuracy threshold (blue dashed line) and HF/CCSD reference lines. The circuit complexity increases with molecular size, requiring longer sequences for more complex molecules. Note that for H2 and LiH, CCSD is extremely close to FCI, so it is not plotted here. Additionally, the kink in N2 around 1.75 Å is due to a sign change in the error.
Figure 6: Comparison of GRPO and Logit Matching loss functions for N2 at bond length 1.5 Å. The plot shows the best energy relative to the ground truth energy as a function of the number of gates (circuit depth).
@@ 3 Results
In this section, we investigate the effectiveness of training in GPT-QE for approximating ground states using electronic structure Hamiltonians. We use the molecular Hamiltonians of $\texttt{H}_{2}$, LiH, $\texttt{BeH}_{2}$, and $\texttt{N}_{2}$ in the sto-3g basis for this purpose. The more detailed process to generate molecular systems is described in Appendix A.1.
The configuration of the GPT-QE model is as follows. Our operator pool consists of Pauli time evolution operators derived from chemically inspired operators based on unitary coupled-cluster single and double excitations (UCCSD), with detailed configuration provided in Appendix A.2. Regarding the transformer model, we employ a configuration based on GPT-2 [7], modified to feature 6 attention layers and 6 attention heads. The hyper-parameter $\beta$ is updated in each epoch of training using the strategy described in Appendix A.3. The initial state, $\rho_{\textrm{0}}$, is set to the Hartree-Fock state.
@@ 3.1 Training Performance
For the training experiments, we set the replay buffer size to $N_{\text{buffer}}^{\textrm{max}}=1,000$, the batch size to $N_{\text{batch}}=50$, the number of circuits generated per epoch to $N_{\text{sample}}=50$, and the number of iterations per epoch to $N_{\text{iter}}=5$. Before starting the training process, we first collected 200 samples to populate the replay buffer, ensuring a sufficient initial dataset for stable training. The training was conducted for different numbers of epochs depending on the molecular complexity: $\texttt{H}_{2}$ was trained for 200 epochs, LiH for 1,000 epochs, and $\texttt{BeH}_{2}$ and $\texttt{N}_{2}$ for 1,500 epochs. For the loss function, we use Group Relative Policy Optimization (GRPO) [26] loss described in Appendix B. All results presented in this section represent the average (standard deviation) over 3 independent trials.
Figure 4 presents the minimum energies found during the training process for four molecular systems ($\texttt{H}_{2}$, LiH, $\texttt{BeH}_{2}$, and $\texttt{N}_{2}$), showing the energy as a function of bond length. The GQE results are displayed as green points, representing the lowest energy values discovered throughout the training epochs, while reference methods are shown as lines: Hartree-Fock (HF) energy in gray dotted lines, coupled cluster singles and doubles (CCSD) energy in blue dashed lines, and full configuration interaction (FCI) energy in black solid lines. The results demonstrate that GQE successfully captures the qualitative behavior of the potential energy surfaces across different molecular systems, with the generated quantum circuits providing reasonable approximations to the ground state energies.
Figure 5 provides a detailed analysis of the absolute error from FCI on a logarithmic scale for each molecule. The blue dashed horizontal line represents the chemical accuracy threshold (1.6 mHa), which is a standard benchmark for quantum chemistry calculations. The plots also include reference lines for HF and CCSD methods to provide context for the GQE performance. The detailed error analysis reveals important insights about the performance of GPT-QE across different molecular systems and bond lengths. For $\texttt{H}_{2}$ and LiH, chemical accuracy (1.6 mHa) is achieved across all target bond lengths, demonstrating the effectiveness of GPT-QE for these relatively simple molecular systems. However, for $\texttt{BeH}_{2}$ and $\texttt{N}_{2}$, the results show a more nuanced behavior: chemical accuracy is maintained near the equilibrium geometry, but deteriorates in the bond dissociation region where the electronic structure becomes more complex. Notably, for $\texttt{N}_{2}$, GPT-QE achieves higher accuracy than CCSD across most bond lengths, particularly in the equilibrium region, highlighting the potential of the generative approach to capture complex electronic correlations that are challenging for traditional quantum chemistry methods.
To further investigate the effectiveness of different loss functions in GPT-QE, we conducted a comparative study between the GRPO and Logit Matching loss functions introduced in Appendix B for the $\texttt{N}_{2}$ molecule at bond length 1.5 Å. Figure 6 shows the best energy relative to the ground truth energy as a function of the number of gates (circuit depth). GRPO consistently outperforms Logit Matching across all gate counts, achieving lower energy values and demonstrating better convergence properties. The chemical accuracy threshold (1.6 mHa) is indicated by the blue dashed line, with the shaded region representing the chemical accuracy regime. This comparison highlights the superior performance of GRPO loss in guiding the generative model towards more accurate quantum circuit generation across different circuit depths.